OpenStudy (anonymous):

Consider the following linear equation: 1/8(z+3)=4/5(z+1/8) Solve the above linear equation. Simplify your answer.

OpenStudy (anonymous):

I've gotcha! ..just let me type it all out.

OpenStudy (dumbcow):

if you want to get rid of the fractions, you can multiply the equation by 40 -> 5(z+3) = 32(z+ 1/8)

OpenStudy (anonymous):

Okay, I hope you can read this: First, simplify the equations: \[1/8(z) + 3/8 = 4/5(z) + 4/40\] Next, collect like-terms (z's) on one side of the equation: \[1/8(z) - 4/5(z) = 4/40 - 3/8\] Multiply through to get a common denominator (of 40): \[5/40(z) - 32/40(z) = 4/40 - 15/40\] Simplify each side by subtracting the like terms on each side: \[-27/40z = -11/40\] Divide the -27/40 into -11/40, or multiply by the reciprocal: \[z = (-11/40) * (40/-27)\] ..and there you have it. z = -11/-27 or just 11/27.